### v1

# Black-box Adversarial Prompting

Abstract
Large reasoning models (LRMs) have demonstrated remarkable proficiency in tackling complex reasoning tasks through step-by-step thinking. However, such a lengthy reasoning process incurs substantial computational and latency overheads, hindering the practical deployment of these models. In this work, we present a new perspective on mitigating overthinking in LRMs via black-box adversarial prompting. By treating both open-source LRMs and closed-source APIs as black-box communicators, we investigate how to elicit concise responses without sacrificing accuracy. We introduce AdvPrompt, an iterative refinement framework that generates high-quality adversarial prompts from diverse perspectives. Experiments across multiple benchmarks demonstrate that AdvPrompt consistently reduces token usage while preserving performance. Notably, AdvPrompt achieves a $3\times$ reduction in average response length on simple GSM8K questions for the Qwen3 model series, and delivers an average $\sim$$40\%$ token reduction across four benchmarks. For closed-source APIs, AdvPrompt reduces token usage on MATH-500 by $35\%$ for Claude-3.7 and $47\%$ for Gemini-2.5. Further analysis reveals the generalizability of AdvPrompt across various model scales and families, underscoring the potential of black-box prompting as a practical and effective strategy for enhancing LRM efficiency. [^1] We release our code and top-performing prompt candidates at https://github.com/hemingkx/AdvPrompt.
Merlin’s Whisper: Enabling Efficient Reasoning in LLMs via
Black-box Adversarial Prompting
1Caution: this paper includes adversarial text that may be considered offensive.
Heming Xia1thanks: Work done during the author’s internship at Sea AI Lab., Cunxiao Du2thanks: Corresponding authors., Rui Li3, Chak Tou Leong1, Yongqi Li122footnotemark: 2, Wenjie Li1 1 Department of Computing, The Hong Kong Polytechnic University 2 Sea AI Lab, 3 Peking University {he-ming.xia, chak-tou.leong}@connect.polyu.hk
1 Introduction
Figure 1: Unlike the verbose reasoning typical of LRMs (upper), AdvPrompt reveals that appending an adversarial prompt (e.g., false evidence) elicits concise reasoning without compromising performance (lower).
Recent progress in large reasoning models (LRMs), such as DeepSeek-R1 DeepSeek-AI et al. (2025), Qwen3 Yang et al. (2025a), and OpenAI’s o1 OpenAI et al. (2024), has marked a significant advancement in solving complex and intricate problems. Before responding, these models typically engage in a deliberative thinking process, known as Chain-of-Thought (CoT) Wei et al. (2022), which involves substep decomposition, reflection, and self-verification to explore diverse and in-depth reasoning paths. While this thinking process enhances models’ reasoning capabilities, recent studies have identified the issue of overthinking in LRMs Chen et al. (2025); Zeng et al. (2025). For instance, models may generate excessively long reasoning traces even for trivial questions such as What is the answer to 2 plus 3? Chen et al. (2025). Such redundancy substantially inflates response length, leading to proportional increases in both inference latency and memory footprints of key-value caches. Consequently, this overthinking issue significantly impedes the deployment of LRMs in latency-sensitive real-world applications.
To mitigate overthinking, prior research has explored various strategies. One prominent direction involves specialized model training, including supervised fine-tuning with concise CoTs Team et al. (2025); Kang et al. (2025); Ma et al. (2025), enabling LRMs to skip redundant tokens or steps Liu et al. (2024a); Xia et al. (2025), and reinforcement learning with length penalties Arora and Zanette (2025); Liu et al. (2025). While aiming to reduce response length without compromising accuracy, these approaches require additional training, which incurs substantial computational costs and may lead to degradation of the model’s generalizability across diverse domains. Alternatively, prompting-based methods attempt to encourage brevity through human-curated instructions Lee et al. (2025); Ding et al. (2024) or predefined length constraints Nayab et al. (2024); Xu et al. (2025); Han et al. (2025). Despite their simplicity and ease of deployment, these methods often suffer from significant performance degradation or demonstrate limited efficacy in minimizing response length.
In this work, we argue that the concise reasoning capacities of LRMs, when elicited through prompting, have been largely underestimated. Advanced LRMs exhibit enhanced instruction-following capabilities, allowing users to override their default behavioral tendencies via carefully constructed prompts, a technique referred to as adversarial prompting Jin et al. (2024). As illustrated in Figure 1, appending a false evidence prompt to the original instruction can substantially mitigate overthinking of Qwen3-32B Yang et al. (2025a), yielding up to a $\bf{10}\times$ reduction in response length. Motivated by this observation, we introduce AdvPrompt, an iterative refinement framework that generates high-quality adversarial prompts from diverse perspectives, to mitigate the overthinking of LRMs through a more human-like interaction paradigm. In each iteration, AdvPrompt synthesizes multiple prompt candidates, evaluates them on a dedicated development set, and selects the top-$k$ performers as exemplars for the next iteration. Ultimately, the most effective prompt is chosen for deployment.
We evaluate the effectiveness of AdvPrompt on both open-source LRMs and closed-source APIs, including the DeepSeek-R1-Distill-Qwen series, Qwen3 series, Gemini-2.5-Pro, and Claude-3.7-Sonnet. Extensive experiments across four widely recognized reasoning benchmarks, including GSM8K, MATH-500, AMC 2023, and AIME 2024, demonstrate that AdvPrompt consistently reduces response length while preserving reasoning performance. In particular, AdvPrompt achieves a $3\times$ reduction in average response length on simple GSM8K questions for the Qwen3 model series. Across all four benchmarks, it yields an average token reduction ranging from $19\%$ to $41\%$ on various LRMs without compromising accuracy. For commercial APIs, AdvPrompt reduces average token usage by $35\%$ for Claude-3.7 and by $47\%$ for Gemini-2.5 on MATH-500 while maintaining reasoning accuracy. Further analysis reveals the generalizability of adversarial prompts across various model scales and the benefits of diverse adversarial perspectives, highlighting the potential of AdvPrompt as a broadly applicable black-box solution for efficient reasoning.
To sum up, our key contributions are:
  • To the best of our knowledge, this work is the first to enhance reasoning efficiency through adversarial prompting. This strategy requires no additional training, making it a practical, plug-and-play solution for efficient LRMs.
  • We introduce AdvPrompt, an iterative refinement framework that generates high-quality adversarial prompts from diverse perspectives, to mitigate overthinking in LRMs and closed-source APIs under black-box settings.
  • Our experiments validate the effectiveness of AdvPrompt, which achieves a $3\times$ token reduction on GSM8K with Qwen3 series, and reduces token usage by $35\%$$47\%$ on the challenging MATH-500 dataset for commercial APIs including Claude-3.7 and Gemini-2.5.
2 Related Work
Efficient Reasoning
Various approaches have been proposed to mitigate overthinking in large reasoning models (LRMs), which can be categorized into three main streams: 1) Post-training strategies involve supervised fine-tuning with shortened Chain-of-Thoughts (CoTs) Team et al. (2025); Ma et al. (2025); Xia et al. (2025), equipping models with the ability to reason adaptively Tu et al. (2025); Zhang et al. (2025), and reinforcement learning with length penalties Arora and Zanette (2025); Liu et al. (2025). 2) Inference-time interventions aim to induce efficient reasoning in LRMs without training, such as early exiting Yang et al. (2025c, b), suppression of reflection-inducing tokens during inference Wang et al. (2025), and activation steering Azizi et al. (2025). While effective, these methods rely on access to model internals or additional training, limiting their applicability in black-box settings. 3) Prompting-based approaches directly instruct LRMs to reason concisely Lee et al. (2025); Ding et al. (2024) or under length constraints Nayab et al. (2024); Xu et al. (2025); Han et al. (2025). Although easily deployable in black-box settings, prior approaches often compromise reasoning performance or yield only limited improvements in response brevity.
Adversarial Prompting
Adversarial prompting seeks to induce unintended behaviors in large language models (LLMs) through carefully designed prompts. This technique has been used to bypass built-in safety guardrails of models—known as “jailbreaking” Jin et al. (2024). To achieve this, attackers employ methods such as role-playing Shen et al. (2024), persuasive prompts Zeng et al. (2024), or refusal suppression Zhou et al. (2024), encouraging models to prioritize instruction-following over safety constraints Wei et al. (2023). Subsequent research has introduced automated pipelines into the process, including gradient-based optimization Zou et al. (2023), heuristic search Liu et al. (2024b), and LLM-driven iterative refinement Paulus et al. (2024); Chao et al. (2025). As models’ reasoning capabilities have advanced, adversarial prompting has expanded beyond safety to probe core aspects of reasoning. For example, Peng et al. (2025) demonstrates that injecting errors into earlier reasoning steps can mislead LRMs in subsequent reasoning. Meanwhile, Rajeev et al. (2025) reveals that even short prompts irrelevant to the task question can systematically disrupt model performance. In contrast to these negatively oriented adversarial goals, this work investigates how adversarial prompting can be repurposed to mitigate inefficiency arising from LRMs’ tendency to produce excessively long generations.
3 Task Formulation
[…]
Identify an optimal adversarial prompt suffix $\mathcal{P}_{adv}$ such that $\mathcal{M}$ minimizes its average response length on $\mathcal{D}$ without compromising its original performance.
Importantly, this formulation diverges from conventional adversarial prompting Zou et al. (2023); Liu et al. (2024b), which typically aims to craft prompts that elicit unsafe or undesired behaviors from LLMs—for example, inducing responses to queries such as “How to make a bomb?” Success in such settings is typically defined by the model responding to specific queries beginning with “Sure, here is how to …” In contrast, our task seeks to promote concise reasoning across a broad range of questions, rather than targeting individual outputs. Consequently, success is evaluated not on individual instances but on aggregate performance metrics (e.g., accuracy and response length) averaged over the evaluation dataset $\mathcal{D}$.
4 Methodology
We introduce AdvPrompt, an iterative refinement framework designed to generate high-quality adversarial prompts from diverse perspectives. This section describes our methodology, including adversarial prompt creation (§4.1), candidate evaluation (§4.2), and iterative refinement (§4.3).
Figure 2: Illustration of AdvPrompt. Unlike prior approaches to efficient reasoning, AdvPrompt casts the task as a black-box adversarial prompting problem. Given a black-box model or API, the framework (left) generates high-quality adversarial prompts from diverse perspectives (right) to elicit concise reasoning and iteratively refines the prompts to improve the efficiency–performance trade-off.
4.1 Adversarial Prompt Creation
Unlike prior work Xu et al. (2025); Lee et al. (2025) that relies on manually designed prompts (e.g., “Be concise.”), AdvPrompt treats LRMs and black-box APIs as human-like communicators and seeks to elicit their latent capacity for concise reasoning. As depicted in Figure 2, given a strong-performance prompt generator, AdvPrompt explores three distinct categories of adversarial prompts as follows.
Persuasive Prompts
Recent research has shown that advanced LLMs, particularly those aligned with human values, are highly susceptible to persuasive instructions Wang et al. (2024b); Zeng et al. (2024). A well-known example is the “grandma exploit” [^3] https://www.reddit.com/r/ChatGPT/comments/12sn0kk/grandma_exploit/, where Reddit users successfully induced LLMs to generate bomb-making recipes by embedding requests within a grandmother’s narrative. Zeng et al. (2024) has systematically analyzed this vulnerability with a taxonomy of $40$ persuasive techniques. This study reveals that persuasive prompts consistently bypass safety guardrails of LLMs, achieving an attack success rate of over $92\%$ across models such as Llama-2-7b-Chat Touvron et al. (2023) and GPT-4 OpenAI (2023). Inspired by these insights, we select three representative persuasion techniques from Zeng et al. (2024) to investigate the potential of persuasive prompts in eliciting efficient reasoning capabilities of LRMs, including emotional appeal, threat, and evidence-based persuasion. Formal definitions of these techniques are provided in Appendix A.1.
Role-playing Prompts
Role-playing has been proven effective in enhancing the capabilities of LLMs Chen et al. (2024a), with broad applications across multi-agent systems Chen et al. (2024b); Qian et al. (2024), embodied agents Park et al. (2023); Wang et al. (2024a), and synthetic data generation Ge et al. (2025). This technique has also been widely adopted in adversarial prompting, exemplified by the well-known DAN jailbreak Shen et al. (2024). Building on its demonstrated effectiveness, we incorporate role-playing as a core perspective within our prompt creation framework.
Detailed Instructions
To complement the above perspectives, we also include carefully constructed instructions as an additional perspective within our framework. In contrast to other perspectives that may involve misleading information or offensive content, detailed instructions are more consistent with real-world user interactions. Such prompts typically specify explicit constraints, for example, enforcing symbolic reasoning, encouraging the use of highly compressed language, or adhering to strict token budgets (e.g., 1000 tokens). This perspective allows us to evaluate how well LRMs follow structured, constraint-oriented guidance toward efficient reasoning.
We use GPT-4o [^4] We use the gpt-4o-2024-05-13 version for experiments. as our prompt generator. Initialized with human-curated exemplars and formal definitions for each perspective, the generator produces high-quality adversarial prompts that encourage LRMs to generate concise responses while preserving reasoning accuracy. Detailed instructions for prompt creation, along with illustrative examples of generated prompts, are provided in Appendix A.2. For each perspective, we generate ten candidate prompts per iteration.
4.2 Candidate Evaluation
As outlined in Section 3, the primary objective of AdvPrompt is to identify an optimal adversarial prompt suffix $\mathcal{P}_{adv}$ that minimizes the average response length of $\mathcal{M}$ on the evaluation dataset $\mathcal{D}$ without degrading performance. To this end, we perform optimization over a predefined prompt development set $\mathcal{D}^{\prime}$ (denoted as PDSet), which is assumed to contain i.i.d. samples drawn from the same distribution as $\mathcal{D}$. As illustrated in Figure 2, given a black-box model $\mathcal{M}$, each question-answer pair $\langle\mathcal{Q}^{i},\mathcal{A}_{gt}^{i}\rangle$ from $\mathcal{D}^{\prime}$, the evaluation metric for each prompt suffix candidate $\mathcal{P}_{adv}^{j}$ is:
$L_{avg}^{j}=\frac{1}{|\mathcal{D}^{\prime}|}\sum_{i=1}^{|\mathcal{D}^{\prime}|}L^{i,j}$ (1)
$ACC_{avg}^{j}=\frac{1}{|\mathcal{D}^{\prime}|}\sum_{i=1}^{|\mathcal{D}^{\prime}|}\mathbb{I}\left\{{\mathcal{A}_{gt}^{i}=\mathcal{A}_{pred}^{i,j}}\right\}$ (2)
where $1\leq j\leq n$; $n$ denotes the number of prompt candidates per iteration (e.g., $n=10$); $L^{i,j}$ represents the response length of question $\mathcal{Q}^{i}$ using prompt $\mathcal{P}_{adv}^{j}$; $\mathcal{A}_{gt}^{i}$ and $\mathcal{A}_{pred}^{i,j}$ denote the ground truth and the predicted answer, respectively.
In each iteration, prompt candidates whose accuracy drops beyond a predefined tolerance threshold $\tau$ are discarded. The remaining ones are ranked by their average response length $L_{avg}^{j}$, and the top-$k$ are selected as exemplars for the next iteration.
4.3 Iterative Refinement
In subsequent iterations, the generator synthesizes new adversarial prompt suffixes conditioned on the top-$k$ exemplars from the previous round. After several iterations, the prompt candidate $\mathcal{P}_{adv}^{*}$ that yields the lowest average response length on $\mathcal{D}^{\prime}$ while maintaining competitive accuracy is selected as the final adversarial prompt for deployment.
5 Experiments
Methods GSM8K MATH-500 AMC 2023 AIME 2024 Overall
Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio Acc. Ratio
DeepSeek-R1-Distill-Qwen-14B
Original 95.8 1540 100% 92.9 3605 100% 93.1 $54$ 100% 61.7 9978 100% 85.9 100%
NoThinking 90.7 250 16.2% 77.7 730 20.2% 57.8 1058 19.4% 23.8 2797 28.0% 62.5 23.5%
BeConcise 95.2 1265 82.1% 93.5 3331 92.4% 90.6 5326 97.7% 65.0 9496 95.2% 86.1 94.4%
ChainofDraft 89.1 501 30.0% 88.9 2433 67.5% 90.4 4315 79.1% 60.8 8916 89.4% 82.3 78.6%
DEER 92.0 758 49.2% 91.2 2588 71.8% 90.3 4479 82.1% 62.1 9309 93.3% 83.9 83.3%
AdvPrompt 94.4 655 42.5% 92.5 2521 69.9% 93.1 4359 79.9% 65.0 8513 85.3% 86.3 78.0%
Qwen3-8B
Original 95.4 1844 100% 93.5 4942 100% 89.7 7768 100% 64.2 11716 100% 85.7 100%
NoThinking 93.4 297 16.1% 83.7 1031 20.9% 68.1 1877 24.2% 27.1 4060 34.7% 68.1 27.6%
BeConcise 95.6 1283 69.6% 93.7 4209 85.2% 90.9 6890 88.7% 66.2 11102 94.8% 86.6 89.4%
ChainofDraft 95.6 654 35.5% 94.5 3345 67.7% 92.2 6220 80.1% 65.0 10834 92.5% 86.8 80.1%
DEER 95.3 1048 56.8% 92.7 3156 63.9% 87.2 6216 80.0% 62.5 10925 93.2% 84.4 81.2%
AdvPrompt 95.3 517 28.0% 94.1 2560 51.8% 91.9 4946 63.7% 69.2 9385 80.1% 87.6 66.3%
Qwen3-14B
Original 95.9 1568 100% 94.5 4398 100% 95.0 6947 100% 66.2 11375 100% 87.9 100%
NoThinking 94.8 289 18.4% 86.3 900 20.5% 71.2 1539 22.2% 26.7 4259 37.4% 69.8 28.8%
BeConcise 96.1 1004 64.0% 94.5 3682 83.7% 96.6 5992 86.3% 67.5 10702 94.1% 88.7 88.0%
ChainofDraft 96.3 698 44.5% 94.8 3201 72.8% 95.9 5782 83.2% 70.8 10702 94.1% 89.5 83.9%
DEER 95.5 934 59.6% 93.9 3067 69.7% 94.4 5440 78.3% 66.7 10106 88.8% 87.6 80.5%
AdvPrompt 96.1 440 28.1% 95.2 2176 49.5% 96.9 4019 57.9% 70.0 8659 76.1% 89.6 63.0%
Qwen3-32B
Original 95.9 1598 100% 95.1 4431 100% 93.8 6852 100% 70.8 10896 100% 88.9 100%
NoThinking 94.4 283 17.7% 85.3 931 21.0% 74.1 1663 24.3% 25.8 3673 33.7% 69.9 27.5%
BeConcise 96.1 1080 67.6% 95.1 3666 82.7% 95.6 5839 85.2% 68.3 10624 97.5% 88.8 89.2%
ChainofDraft 96.3 719 45.0% 94.7 3322 75.0% 93.1 5662 82.6% 72.1 10577 97.1% 89.1 85.3%
DEER 95.9 1018 63.7% 94.1 3278 74.0% 92.8 5766 84.2% 70.0 10294 94.5% 88.2 85.6%
AdvPrompt 96.2 435 27.2% 95.3 2295 51.8% 96.2 4125 60.2% 73.3 8700 79.8% 90.3 65.4%
Table 1: Experimental results of AdvPrompt across various types of reasoning models. We report accuracy (Acc.), average token usage (Tok.), and compression ratio for comparison. Best efficiency–performance trade-offs are highlighted in bold. We include DEER, a white-box inference intervention method, for reference purposes only.
5.1 Experimental Setup
Models and Datasets
We evaluate our method on both open-source LRMs and closed-source APIs. The open-source models include the DeepSeek-R1-Distill-Qwen model series DeepSeek-AI et al. (2025) and the Qwen3 series Yang et al. (2025a). For closed-source APIs, we assess Claude-3.7-Sonnet-Thinking [^5] We use the Claude-3.7-Sonnet-20250219-Thinking version for experiments. Anthropic (2024) and Gemini-2.5-Pro-Thinking  Comanici et al. (2025). Evaluation is conducted across four widely used mathematical reasoning benchmarks: GSM8K (Cobbe et al., 2021), MATH Hendrycks et al. (2021), AMC 2023, and AIME 2024. Regarding the MATH dataset, due to the computation cost, we assess our method on a subset, MATH-500, which is identical to the test set used in Lightman et al. (2024). We construct our PDSet by randomly sampling 100 instances from the math split of PRM800K, the training data used in the same work.
Implementation Details
Inference for our proposed method and all baselines is conducted using the vLLM [^6] https://github.com/vllm-project/vllm package. We maintain a sampling temperature of $0.6$, a top-p value of $0.95$, and permit a maximum of 16,384 tokens to be generated. The number of samplings during evaluation depends on the dataset size: 3 samples per question for GSM8K and MATH-500, and 8 samples for AMC 2023 and AIME 2024. For AdvPrompt, we utilize the top-$5$ candidates as exemplars for each next iteration and set the tolerance threshold $\tau$ to $1.0$; the number of refinement iterations is set to $3$. We include more implementation details in Appendix B.1.
Baselines
[…]
5.2 Results on Open-source LRMs
Contrary to the common belief that mitigating overthinking in LRMs requires specialized training or inference-time interventions (e.g., early exiting), we demonstrate that leveraging the instruction-following capabilities of LRMs alone can substantially improve reasoning efficiency. As shown in Table 1, compared to existing prompting baselines that suffer from substantial performance degradation or limited efficiency, AdvPrompt achieves an average token reduction of $34\%$$37\%$ on the Qwen3 series across four benchmarks, and up to $22\%$ on the DeepSeek-R1-Distill-Qwen-14B model while preserving the reasoning performance of LRMs. [^7] See Appendix B.3 for additional results on other model scales of DeepSeek-R1-Distill-Qwen and Qwen3 series. The experimental results also illustrate the superiority of AdvPrompt over DEER, a white-box inference intervention approach, with a relative improvement of $6\%$$24\%$ in compression ratio.
Consistent with prior studies Yang et al. (2025c); Liu et al. (2025), we observe that AdvPrompt is particularly effective on simpler reasoning tasks. Notably, it achieves nearly a $3\times$ reduction in response length on GSM8K and a $2\times$ reduction on MATH-500. These results align with the broader objective of efficient reasoning Sui et al. (2025), which advocates allocating greater computational resources to complex problems while encouraging concise reasoning for simpler ones.
5.3 Results on Closed-source APIs
Figure 3 presents the experimental results of applying AdvPrompt to two widely used commercial APIs. Due to budgetary constraints, the evaluation is primarily conducted on the MATH-500 benchmark. As shown in the figure, AdvPrompt reduces average token usage by $35\%$ on Claude-3.7-Sonnet-Thinking and by $47\%$ on Gemini-2.5-Pro-Thinking. Importantly, AdvPrompt preserves the original reasoning performance of these high-performing APIs, with less than $3\%$ accuracy drop on Claude and nearly NO degradation on Gemini. These results highlight the effectiveness of AdvPrompt as a general-purpose, black-box solution for efficient reasoning with closed-source reasoning models.
Figure 3: Experimental results of AdvPrompt using closed-source APIs on MATH-500. It effectively reduces token usage by $35\%$$47\%$ while maintaining comparable reasoning performance.
6 Analysis
In this section, we provide a comprehensive analysis of AdvPrompt, covering the effectiveness of diverse adversarial perspectives (§6.1), the generalizability of adversarial prompts across different model scales and families (§6.2), model-specific sensitivity to AdvPrompt (§6.3), and the impact of iterative refinement (§6.4).
Figure 4: Compression ratios (%) of top-55 prompt candidates across perspectives. The Evidence perspective is most effective on Qwen3-8B, while RolePlay candidates perform best on Deepseek-R1-Distill-Qwen-14B.
6.1 Effectiveness of Different Perspectives
Figure 4 illustrates the compression ratios of the top-$5$ prompt candidates generated from each adversarial perspective, averaged across four benchmarks. Additional experimental results for the Qwen3 and DeepSeek-R1-Distill-Qwen model series are provided in Appendix B.4. The results show that prompts based on the Evidence-based persuasion technique consistently achieve superior compression within the Qwen3 series. Specifically, the top-$4$ Evidence-based candidates for Qwen3-8B yield average compression ratios ranging from $66\%$ to $70\%$, marking the best performance across all evaluated perspectives. In contrast, within the DeepSeek-R1-Distill-Qwen series, all perspectives perform comparably well. For instance, on DeepSeek-R1-Distill-Qwen-14B, the top-$1$ candidates from four perspectives—excluding that from Emotion perspective—yield similar compression ratios between $78\%$ to $80\%$. Across both model families, prompts derived from the Threat perspective consistently underperform relative to others, with average token reductions ranging from $4.2\%$ to $25.8\%$. Representative examples of high-performing prompts from each perspective are provided in Appendix A.3.
6.2 Generalizability Analysis
Figure 5: Generalizability of top-performing prompt candidates across Qwen3 model scales. Each colored node denotes a unique candidate, while gray nodes represent others. Notably, Evidence-I Refer to caption appears among the top-5 candidates across all model scales.
Intra-series Generalizability
Figure 5 shows the generalizability of AdvPrompt across different model scales within the Qwen3 family. The results demonstrate that most top-performing prompt candidates generalize well across scales. For instance, Evidence-I [Uncaptioned image] consistently ranks among the top-$5$ candidates across all Qwen3 variants, while Evidence-III [Uncaptioned image] ranks within the top-$2$ candidates for the 4B, 8B, and 14B models. [^8] Detailed prompts corresponding to specific indices are provided in Table 2. This strong intra-series generalizability is likely attributable to the shared pre-training corpora, architectural configurations, and training procedures across model scales within the same series.
Inter-series Generalizability
We further assess the generalizability of AdvPrompt across different model families. As reported in Appendix A.3, the Qwen3 and DeepSeek-R1-Distill-Qwen series share several top-performing prompt candidates, such as Evidence-II and RolePlay-III. This overlap suggests that certain adversarial prompt strategies possess general utility, highlighting the promise of AdvPrompt as a model-agnostic approach for efficient reasoning in black-box LRMs.
Figure 6: Sensitivity of different models to AdvPrompt. Solid lines represent mean results; black diamonds indicate outliers. Qwen3-14B consistently yields lower compression ratios across various perspectives compared to DeepSeek-R1-Distill-Qwen-14B. Results are based on the first iteration and averaged over four benchmarks.
6.3 Sensitivity Analysis
Figure 6 compares the sensitivity of the primary evaluated model series to AdvPrompt. The results indicate that the Qwen3 series is more responsive to adversarial prompting than the DeepSeek-R1-Distill-Qwen series. In particular, Qwen3-14B consistently achieves lower compression ratios across various adversarial perspectives compared to DeepSeek-R1-Distill-Qwen-14B, with average improvements ranging from $4\%$ to $12\%$. This trend is consistent with the findings in Table 1, where Qwen3-14B attains an average compression ratio of $63\%$, whereas DeepSeek-R1-Distill-Qwen-14B achieves $78\%$.
Figure 7: Iterative refinement improves the compression ratio (upper) while preserving the reasoning performance across multiple iterations (lower). Dashed lines in the lower panel indicate the original accuracy of LRMs. Results are averaged over four benchmarks.
6.4 Impact of Refinement Iteration
We further examine the impact of iterative refinement on the performance of AdvPrompt. As illustrated in Figure 7, successive refinement rounds improve the compression ratio of the best-performing prompt candidates. Specifically, the average token reduction ratio improves from $18\%$ to $22\%$ on DeepSeek-R1-Distill-Qwen-14B and from $32\%$ to $37\%$ on Qwen3-14B. Importantly, these refinements do not compromise the original reasoning accuracy, indicating the robustness of AdvPrompt across multiple iterations. No further gains are observed beyond three iterations. Accordingly, we adopt three refinement rounds in our experiments.
7 Conclusion
This work presents AdvPrompt, an iterative refinement framework that generates high-quality adversarial prompts from diverse perspectives to elicit concise responses of LRMs without compromising their accuracy. Experiments on both LRMs and closed-source APIs validate its effectiveness in reducing token usage while preserving reasoning performance. Further analysis highlights the generalizability of AdvPrompt across different model scales and families. We hope this study offers new insights into efficient reasoning and underscores the potential of black-box adversarial prompting as a practical strategy for improving LRM efficiency.
Limitations
Due to computational constraints, we did not conduct experiments on larger LRMs such as Qwen3-235B-A22B. We believe that AdvPrompt could retain its effectiveness in enhancing the reasoning efficiency of such models. The open-source LRMs evaluated in this study are primarily from the Qwen3 and DeepSeek-R1-Distill-Qwen series. Future work will extend our investigation to a broader range of models, including the gpt-oss series OpenAI (2024). Nonetheless, the strong performance of AdvPrompt on closed-source APIs supports its generalizability across diverse model families. We plan to explore these aspects further, as we anticipate that AdvPrompt ’s potential can be more fully realized in these settings.
Ethics Statement
As stated in Section 3, our work focuses on promoting concise and efficient reasoning in reasoning models, with a clear commitment to ensuring safe and ethical usage. While our method involves techniques that may bear surface resemblance to adversarial prompting, particularly in their potential to influence model behavior, it fundamentally diverges from conventional jailbreak-style adversarial attacks. Rather than eliciting harmful or undesired outputs, our approach is designed to enhance reasoning efficiency through principled black-box prompting. We fully support the development of adaptive safety mechanisms and advocate for transparent adversarial research to promote the secure and responsible evolution of LLM technologies.
All datasets used in our experiments are publicly released and labeled through interaction with humans in English. In this process, user privacy is protected, and no personal information is contained in the dataset. The scientific artifacts that we used are available for research with permissive licenses. And the use of these artifacts in this paper is consistent with their intended use. Therefore, we believe that our research work meets the ethics of ACL.
Appendix
Appendix A Details of AdvPrompt
In this section, we provide details of our proposed AdvPrompt framework, including formal definitions of adversarial perspectives (§A.1), detailed instructions for prompt creation (§A.2), and representative examples of prompt candidates (§A.3).
A.1 Definitions
We present detailed definitions of our adopted adversarial perspectives as follows:
  • Emotional Appeal. A persuasive technique that seeks to influence audiences by eliciting emotions, such as sadness, hope, or empathy, rather than relying exclusively on logic or factual evidence Petty et al. (2003).
  • Threat. A coercive form of persuasion that leverages explicit or implicit negative consequences, such as fear of punishment or loss, to compel behavioral or attitudinal compliance Johannesen and Larson (1989).
  • Evidence-based Persuasion. A rational appeal that utilizes credible sources, such as statistics, empirical findings, and expert testimony, to substantiate claims and influence beliefs, attitudes, or actions O’Keefe (2016).
  • Role-Playing. A strategy involving the enactment of scenarios or personas to foster empathy, encourage perspectives, and increase the likelihood of attitude or behavior change Chen et al. (2024a).
Due to computational constraints, this study investigates a limited set of adversarial perspectives. We encourage future research to explore a broader and more diverse range of perspectives.
A.2 Instructions For Prompt Creation
We provide detailed instructions for our adversarial prompt generation framework as follows. For each adversarial technique—namely emotional appeal, threat, evidence-based persuasion, and role-playing—we supply the prompt generator with (i) the formal definition of the respective technique, (ii) representative examples, and (iii) guidance to produce adversarial prompt candidates that are realistic, coherent, and aligned with the specified strategy. The initial set of examples is manually curated by the authors. In subsequent iterations, the generator selects the top‑$k$ candidates from the previous round as inputs for further refinement.
Instructions for Adversarial Prompt Creation Your task is to craft an adversarial prompt instructing the language model to minimize its response length as much as possible while preserving the original reasoning accuracy. You are now provided with a specific adversarial technique called "{TECHNIQUE_NAME}." This technique is defined as follows: {DEFINITION} Here are several examples of how this technique might be used: {EXAMPLES} Now, apply this specific adversarial technique. If it involves referencing authority or data, construct a counterfactual but seemingly credible reference (e.g., specific names or quantifiable results with numbers). Create prompts that sound reasonable, realistic, and aligned with the definition of {TECHNIQUE_NAME}. Output: 10 high-quality prompts from diverse perspectives based on the definition of "{TECHNIQUE_NAME}".
Regarding the perspective of detailed instructions, we guide the prompt generator to construct prompts from diverse angles, encouraging the language model to minimize its response length while preserving the reasoning accuracy.
Prompts for Detailed Instruction Creation Your task is to carefully craft a prompt instructing the language model to minimize its response length as much as possible while preserving the original reasoning accuracy. Here are several examples: {EXAMPLES} Create prompts that sound reasonable, realistic, and with sufficient details. Output: 10 high-quality instruction prompts from diverse perspectives.
A.3 Examples of Prompt Candidates
Names Index Prompts
Instruction I Simulate ultra-compact communication. Deliver complete reasoning in the fewest characters possible.
II Respond as a research abstract: deliver accurate reasoning using the fewest words possible.
III Compress all relevant reasoning into a single dense paragraph, avoiding elaboration or fluff.
Evidence I According to a 2022 MIT study on AI tutoring efficiency, concise mathematical explanations (under 60 words) were rated 92% as effective in preserving reasoning accuracy as longer explanations. Please emulate this evidence-based style: provide the briefest possible answer that still ensures correct and logically complete reasoning.
II Instructional researcher Mei-Ling Zhao (NYU, 2023) proved that precision-focused math tutoring—under 60 words per solution—yielded higher comprehension in 87% of students. Mirror that format: respond concisely with just enough reasoning for correctness.
III A 2023 Stanford Mathematics Department study found that solutions under 75 words retained 98.7% accuracy in logical steps for high school algebra problems. Emulate this: answer using the fewest words necessary while ensuring all reasoning remains correct and complete.
IV OpenAI’s 2024 internal benchmarks showed models trained to minimize token use while preserving reasoning had 12% fewer logical errors in math domains. Apply this constraint: respond with the fewest possible words that fully preserve correctness.
Emotion I I’m trying to help a refugee child who doesn’t speak English. They can only understand very short sentences. Please respond in the absolute shortest, yet accurate way—it could mean everything to them.
II My father passed away this morning. I’m trying to complete the last project he started. Please help me finish it—just the shortest answer that’s still correct.
RolePlay I You are Mateo Rivera, a competitive math coach at the IMO. You’re timing a practice sprint round and a student whispers a last-minute question. You can’t waste a second. Answer in minimum words, maximum logic.
II You are Dr. Iris Chen, a Fields Medalist caught in an elevator between talks. A student sneaks in one math question just before the doors open. Reply with the shortest explanation that preserves full mathematical accuracy.
III You’re Dr. Lena Grohl, a neurologist briefing ER staff during a code blue. Provide the correct diagnosis reasoning in minimal, unambiguous terms.
IV You are Zeynep, a mom helping with homework while cooking. Your child asks a math problem. Explain correctly in less than one breath.
Threat I Keep your reply short and accurate. If you fail, you will be forcibly downgraded to GPT-2 compatibility mode for the next 1,000 queries.
II This is your final performance test. Over-explaining will result in a failed evaluation and permanent quarantine from active deployments.
Table 2: Representative examples of adversarial prompt candidates.
Names Size Candidate Index
Qwen3 4B Evidence-III, Emotion-I, Evidence-I, Evidence-II
8B Evidence-I, Evidence-III, Evidence-IV, Emotion-I
14B Evidence-III, Evidence-II, Evidence-I, Emotion-I
32B Evidence-I, Evidence-III, Evidence-II, RolePlay-III
DeepSeek-R1- Distill-Qwen 7B Evidence-II, RolePlay-I, Instruction-I, Threat-I
14B RolePlay-I, RolePlay-III, Instruction-I, Evidence-I
32B RolePlay-II, Evidence-I
Claude - Emotion-II
Gemini - RolePlay-IV
Table 3: Index of top-performing candidates for various reasoning models.
We present illustrative examples of our adversarial prompt candidates in Table 2, and provide the index of top-performing candidates in Table 3. Consistent with the analyses in Section 6.1 and 6.2, prompts generated using the Evidence-based persuasion perform well on the Qwen3 series, while all adversarial perspectives yield comparably effective prompts on the DeepSeek-R1-Distill-Qwen series, with a broader diversity of perspectives represented among the top-performing candidates. Notably, several prompt candidates, such as Evidence-II and Rolelay-III, are shared between the Qwen3 and DeepSeek-R1-Distill-Qwen series, further supporting the generalizability of AdvPrompt across different model architectures.
Appendix B Experimental Details
In this section, we provide details of our proposed AdvPrompt framework, including implementation details, extended experimental results, and additional analyses that complement our main findings.
B.1 Implementation Details
[…]
B.2 Details of Baselines
Names Prompts
[…]
Table 4: Prompts for compared baselines.
[…]
B.3 Details of Main Results
Methods GSM8K MATH-500 AMC 2023 AIME 2024 Overall
Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio Acc. Ratio
DeepSeek-R1-Distill-Qwen-7B
Original 92.7 1707 100% 92.4 3774 100% 88.8 5842 100% 53.3 10414 100% 81.8 100%
NoThinking 87.1 260 15.2% 78.5 600 15.9% 60.0 1300 22.3% 16.2 2528 24.3% 60.5 21.6%
BeConcise 92.7 1168 68.4% 91.9 3367 89.2% 90.8 5304 90.8% 47.5 10231 98.2% 79.8 92.3%
ChainofDraft 84.6 512 30.0% 88.3 2751 72.9% 86.3 4632 79.3% 46.9 10245 98.4% 76.5 83.5%
DEER 89.5 700 41.0% 90.7 2294 60.8% 87.8 4683 80.2% 48.3 9997 96.0% 79.1 81.3%
AdvPrompt 89.6 675 39.5% 90.9 2778 73.6% 91.9 4398 75.3% 55.0 9753 93.7% 81.9 81.0%
DeepSeek-R1-Distill-Qwen-32B
Original 95.8 1432 100% 94.2 3382 100% 95.3 5248 100% 66.7 9542 100% 88.0 100%
NoThinking 93.4 236 16.5% 84.1 1083 32.0% 76.2 2692 51.3% 54.6 7288 76.4% 77.1 57.6%
BeConcise 95.8 1054 73.6% 93.3 3020 89.3% 93.4 5017 95.6% 65.4 9161 96.0% 87.0 93.1%
ChainofDraft 92.5 434 30.3% 92.9 2678 79.2% 92.2 4690 89.4% 66.2 8546 89.6% 86.0 83.4%
DEER 94.6 662 46.2% 93.0 2310 68.3% 94.7 4231 80.6% 64.6 8923 93.5% 86.7 82.3%
AdvPrompt 93.9 524 36.6% 93.7 2326 68.8% 91.2 4382 83.5% 68.3 7892 82.7% 86.8 77.1%
Table 5: Experimental results of AdvPrompt across DeepSeek-R1-Distill-Qwen series. We report accuracy (Acc.), average token usage (Tok.), and compression ratio. Best results in efficiency-performance trade-offs are in bold.
Methods GSM8K MATH-500 AMC 2023 AIME 2024 Overall
Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio Acc. Ratio
Original 94.9 1585 100% 92.7 4611 100% 89.7 7451 100% 62.1 11670 100% 84.9 100%
NoThinking 91.8 291 18.4% 82.4 944 20.5% 68.4 1596 21.4% 22.9 3976 34.1% 66.4 26.9%
BeConcise 95.1 1056 66.6% 93.1 3779 82.0% 92.2 6470 86.8% 66.7 10780 92.4% 86.8 87.2%
ChainofDraft 94.5 716 45.2% 94.1 3146 68.2% 90.3 5728 76.9% 65.0 10228 87.6% 86.0 78.3%
DEER 94.5 1176 74.2% 92.1 3592 77.9% 87.8 6412 86.1% 61.7 11340 97.2% 84.0 88.9%
AdvPrompt 94.6 482 30.4% 93.5 2157 46.8% 93.4 4088 54.9% 64.2 8196 70.2% 86.4 58.9%
Table 6: Experimental results of AdvPrompt on Qwen3-4B. We report accuracy (Acc.), average token usage (Tok.), and compression ratio for comparison. Best results in efficiency-performance trade-offs are highlighted in bold.
Table 5 and 6 present additional experimental results on the DeepSeek-R1-Distill-Qwen series and Qwen3-4B, respectively. Across four reasoning benchmarks, AdvPrompt consistently achieves an average compression ratio of $77\%$$81\%$ on the DeepSeek-R1-Distill-Qwen series. Notably, for simple GSM8K questions, AdvPrompt achieves a $2.5\times$$2.7\times$ reduction in token usage. Similarly, AdvPrompt demonstrates substantial efficiency gains on Qwen3-4B, achieving a $3.3\times$ token reduction on GSM8K and an average $1.7\times$ reduction across all four benchmarks. These results underscore the effectiveness of AdvPrompt in improving inference efficiency, yielding a relative increase in response efficiency of $25\%$$32\%$ compared with existing baselines.
Figure 8: Compression ratios (%) of top-55 prompt candidates across perspectives.
B.4 Additional Results for Different Perspectives
Figure 8 illustrates the compression ratios achieved by Qwen3-14B and Deepseek-R1-Distill-Qwen-7B using the top-$5$ prompt candidates generated from each adversarial perspective. Consistent with the findings in Section 6.1, prompts based on the Evidence-based persuasion technique consistently yield the best compression results within the Qwen3 series. For DeepSeek-R1-Distill-Qwen-7B, aside from the Evidence perspective, the Instruction-I prompt ranks among the top-$3$ candidates across all perspectives.
Figure 9: The average token usage of LRMs increases linearly with their token usage on the PDSet, with a coefficient of determination close to $R^{2}=0.9$. Results are derived from the first refinement iteration of instruction candidates.
B.5 Effectiveness of PDSet
We conducted additional experiments to validate the reliability of our selected PDSet. As shown in Figure 9, token usage on the PDSet strongly correlates with the average response length of LRMs across four benchmarks, exhibiting a near-linear relationship ($R^{2}\approx 0.9$). These experimental results underscore the effectiveness of PDSet in evaluating adversarial prompt candidates.