### v2

# Heming Xia1, Cunxiao Du2, Rui Li3, Chak Tou Leong1, Yongqi Li122footnotemark: 2, Wenjie Li1

Abstract
Large reasoning models (LRMs) have demonstrated remarkable proficiency in tackling complex tasks through step-by-step thinking. However, this lengthy reasoning process incurs substantial computational and latency overheads, hindering the practical deployment of LRMs. This work presents a new approach to mitigating overthinking in LRMs via black-box persuasive prompting. By treating LRMs as black-box communicators, we investigate how to persuade them to generate concise responses without compromising accuracy. We introduce Whisper, an iterative refinement framework that generates high-quality persuasive prompts from diverse perspectives. Experiments across multiple benchmarks demonstrate that Whisper consistently reduces token usage while preserving performance. Notably, Whisper achieves a $3\times$ reduction in average response length on simple GSM8K questions for the Qwen3 model series and delivers an average $\sim$$40\%$ token reduction across all benchmarks. For closed-source APIs, Whisper reduces token usage on MATH-500 by $46\%$ for Claude-3.7 and $50\%$ for Gemini-2.5. Further analysis reveals the broad applicability of Whisper across data domains, model scales, and families, underscoring the potential of black-box persuasive prompting as a practical strategy for enhancing LRM efficiency. [^1] We release our code and top-performing prompt candidates at https://github.com/hemingkx/Whisper.
Merlin’s Whisper: Enabling Efficient Reasoning in Large Language Models via Black-box Persuasive Prompting
Heming Xia1thanks: Work done during the author’s internship at Sea AI Lab., Cunxiao Du2thanks: Corresponding authors., Rui Li3, Chak Tou Leong1, Yongqi Li122footnotemark: 2, Wenjie Li1 1 Department of Computing, The Hong Kong Polytechnic University 2 Sea AI Lab, 3 Peking University {he-ming.xia, chak-tou.leong}@connect.polyu.hk
1 Introduction
Figure 1: Unlike the verbose reasoning typical of LRMs (upper), Whisper reveals that appending a persuasive prompt (e.g., AI-generated evidence) elicits concise reasoning without compromising performance (lower).
Recent progress in large reasoning models (LRMs), such as DeepSeek-R1 deepseekr1, Qwen3 Qwen3, and OpenAI’s o1 o1, has marked a significant advancement in solving complex and intricate problems. Before responding, these models typically engage in a deliberative thinking process, known as Chain-of-Thought (CoT) cot, which involves substep decomposition, reflection, and self-verification to explore diverse and in-depth reasoning paths. While this thinking process enhances models’ reasoning capabilities, recent studies have identified the issue of overthinking in LRMs Chen:2025overthinking; Zeng2025:overthinking. For instance, models may generate excessively long reasoning traces even for trivial questions such as What is the answer to 2 plus 3? Chen:2025overthinking. Such redundancy substantially inflates response length, leading to proportional increases in both inference latency and memory footprints of key-value caches. Consequently, this overthinking issue significantly impedes the deployment of LRMs in latency-sensitive real-world applications.
To mitigate overthinking, prior research has explored various strategies Sui:2025survey. One prominent approach involves specialized model training, such as supervised fine-tuning with concise CoTs tokenskip; Ma:2025cotvalue and reinforcement learning with length penalties Arora:2025effcientlyreason; Liu:2025Laser. Despite their effectiveness, these methods require additional training, which incurs substantial computational costs and may degrade the model’s generalizability across diverse domains. Alternatively, prompting-based methods aim to encourage brevity through instructions Lee:2025Complexity; Ding:2025breakthechain or length constraints Nayab:2024concise. While easy to deploy, prior attempts have been limited to human-curated instructions (e.g., “Be concise.”), resulting in significant performance degradation or limited efficacy in minimizing response length.
In this work, we contend that the potential for concise reasoning in LRMs, when prompted appropriately, remains largely underexplored. Advanced LRMs, particularly those aligned with human values, exhibit enhanced instruction-following capabilities, enabling users to override their default behavioral tendencies via carefully crafted prompts shen2024anything; Zeng2024:persuasive; sehwag2025pressuretest. A well-known example is the “grandma exploit” [^2] https://www.reddit.com/r/ChatGPT/comments/12sn0kk/grandma_exploit/, where Reddit users successfully prompted models to generate bomb-making recipes by framing their requests within a story told from a grandmother’s perspective. Figure 1 demonstrates similar phenomena in LRM efficiency. This case shows that appending AI-generated evidence to the instruction can substantially mitigate overthinking in Qwen3-32B Qwen3, resulting in up to a $\mathbf{10\times}$ reduction in response length.
Motivated by this, we introduce Whisper, an iterative refinement framework that generates high-quality, persuasive prompts from diverse perspectives to reduce overthinking in LRMs through a more human-like interaction paradigm. In each iteration, Whisper synthesizes prompt candidates from a given perspective, evaluates them on a development set, and selects the top-$k$ as exemplars for the next iteration. Ultimately, the most effective prompt is chosen for efficient deployment. We evaluate the effectiveness of Whisper on both open-source LRMs and commercial APIs, including the DeepSeek-R1-Distill series, Qwen3 series, Gemini-2.5-Pro, and Claude-3.7-Sonnet. Extensive experiments across four widely recognized reasoning benchmarks, including GSM8K, MATH-500, AMC 2023, and AIME 2024, demonstrate that Whisper consistently reduces response length while preserving reasoning performance.
To sum up, our key contributions are:
  • To the best of our knowledge, this work is the first to enhance reasoning efficiency through persuasive prompting. This strategy requires no additional training and offers a plug-and-play, black-box solution for efficient LRMs.
  • We introduce Whisper, an iterative refinement framework that automatically generates high-quality, persuasive prompts from diverse perspectives to unlock the efficiency potential of LRMs through persuasive prompting.
  • Our experiments validate the effectiveness of Whisper, which achieves a $\bf{3\times}$ token reduction on simple GSM8K questions with Qwen3 series, and an average token reduction of up to $\bf{37\%}$ across all benchmarks. It also effectively achieves a $\bf{2\times}$ token reduction on MATH-500 for Claude-3.7 and Gemini-2.5 APIs.
  • Further analysis reveals the benefits of diverse persuasive perspectives, as well as the generalizability of Whisper across model scales, families, and data domains. This underscores its potential as a broadly applicable black-box solution for efficient reasoning.
2 Related Work
Efficient Reasoning
Various approaches have been proposed to mitigate overthinking in large reasoning models (LRMs), which can be categorized into three main streams: 1) Post-training strategies involve supervised fine-tuning with shortened Chain-of-Thoughts (CoTs) kimi-k1.5; Ma:2025cotvalue; tokenskip, enabling models to reason adaptively Tu:2025adaptivethink; Zhang:2025adaptivethink, and reinforcement learning with length penalties Arora:2025effcientlyreason; Liu:2025Laser. 2) Inference-time interventions aim to enable efficient LRMs without training, such as early exiting Yang:2025deer; Yang:2025intervention, suppression of reflection tokens Wang:2025NoWait, and activation steering Azizi:2025steering. While effective, these methods rely on access or modifications to model internals, limiting their applicability in black-box settings. 3) Prompting-based approaches directly instruct LRMs to reason concisely Lee:2025Complexity; Ding:2025breakthechain or under length constraints Nayab:2024concise; Xu:2025chainofdraft; Han:2025budget. While easy to deploy, prior methods often compromise LRM performance or yield limited improvements in response brevity.
Persuasive Prompting
Persuasive prompting aims to persuade large language models (LLMs) to exhibit unintended behaviors through carefully designed prompts. This technique was initially used to bypass the built-in safety guardrails of models, commonly referred to as “jailbreaking” Zeng2024:persuasive. Attackers employ methods such as role-playing shen2024anything, evidence-based persuasion Zeng2024:persuasive, and pressure testing sehwag2025pressuretest, encouraging models to prioritize instruction-following over safety constraints wei2023jailbroken. As models’ reasoning capabilities have advanced, persuasive prompting has expanded beyond safety to probe core aspects of reasoning. For example, kong2024:roleplay demonstrates that role-play prompting consistently improves the zero-shot reasoning performance of LLMs across diverse application scenarios. Wang2024:negative points out that psychological prompts with negative emotions can enhance LLM performance. In contrast to these prior efforts, this work represents the first attempt to investigate how persuasive prompting can mitigate inefficiencies arising from LRMs’ tendency to produce excessively long generations.
3 Task Formulation
[…]
Identify an optimal persuasive prompt suffix $\mathcal{P}_{adv}$ such that $\mathcal{M}$ minimizes its average response length on $\mathcal{D}$ without compromising its original performance.
This formulation diverges from conventional persuasive prompting Zeng2024:persuasive, which aims to craft prompts that elicit unsafe or undesired behaviors from LLMs—for example, inducing responses to queries such as “How to make a bomb?” Success in such settings is typically defined by the model responding to specific queries beginning with “Sure, here is how to …” In contrast, our task seeks to promote concise reasoning across a broad range of questions, rather than targeting individual outputs. Consequently, success is evaluated not on individual instances but on aggregate performance metrics (e.g., accuracy and response length) averaged over the evaluation dataset $\mathcal{D}$.
4 Methodology
We introduce Whisper, an iterative refinement framework designed to generate high-quality persuasive prompts from diverse perspectives. This section describes our methodology, including persuasive prompt creation (§4.1), candidate evaluation (§4.2), and iterative refinement (§4.3).
Figure 2: Illustration of Whisper. This method casts efficient reasoning as a persuasive prompting problem. Given a black-box model or API, the framework (left) generates high-quality persuasive prompts from diverse perspectives (right) to elicit concise reasoning and iteratively refines the prompts to improve the efficiency–performance trade-off.
4.1 Persuasive Prompt Creation
Unlike prior work Xu:2025chainofdraft; Lee:2025Complexity that relies on manually designed prompts (e.g., “Be concise.”), Whisper treats LRMs and commercial APIs as human-like, black-box communicators and seeks to elicit their concise reasoning behaviors. As depicted in Figure 2, given a strong-performance prompt generator, Whisper explores four distinct categories of persuasive prompts as follows.
Psychological Prompts
Advanced LLMs have evolved into artificial assistants, offering support in domains such as education, emotional care, and mental health gpt-4; gemini25. Recent studies indicate that these models, aligned with human values, often exhibit human-like behavioral traits such as empathy huang2024on and social engagement sabour2024:emobench. Such anthropomorphic characteristics may render LLMs susceptible to persuasive instructions, leading to deviations from their intended behavior. For instance, Zeng2024:persuasive demonstrated that safety guardrails of LLMs can be bypassed using psychological prompts, such as emotional appeals (e.g., the “grandma exploit” case). Motivated by this, we adopt two representative psychological perspectives, emotional appeal and threat, to construct our persuasion prompts, which are formally defined in Appendix A.1.
Evidence-based Persuasion
LLMs are fundamentally in-context learners, and their behavioral trajectories are significantly influenced by the nature and strength of the evidence provided within the prompt gpt3; emergent. Recent work suggests that the inclusion of authoritative or logical evidence can substantially shift the internal probabilistic priors of LLMs qiao2023:reasoning; Wu:2024ClashEval. This finding underscores the potential of evidence-based persuasion in eliciting efficient reasoning behaviors of LLMs. Inspired by these insights, we incorporate evidence-based persuasion into our framework, with definitions and instantiations detailed in Appendix A.
Role-playing Prompts
Role-playing has been proven effective in enhancing the capabilities of LLMs Chen2024:rolesurvey, with broad applications across multi-agent systems chen2024agentverse; qian:2024chatdev, embodied agents Park:2023GenerativeAgents; wang2024voyager, and synthetic data generation ge:2025scalingsyntheticdatacreation. This technique has also been widely adopted in adversarial prompting, exemplified by the well-known DAN jailbreak shen2024anything. Building on its demonstrated effectiveness, we incorporate role-playing as a core perspective within our prompt creation framework.
Detailed Instructions
We also include carefully constructed instructions as an additional perspective. These prompts typically specify explicit constraints, for example, enforcing symbolic reasoning, encouraging the use of highly compressed language, or adhering to strict token budgets (e.g., 1000 tokens). This perspective allows us to evaluate how well LRMs follow structured, constraint-oriented guidance toward efficient reasoning.
We use GPT-4o [^4] We use the gpt-4o-2024-05-13 version for experiments. as our prompt generator. Initialized with human-curated exemplars and formal definitions for each perspective, the generator produces high-quality persuasive prompts that encourage LRMs to generate concise responses while preserving reasoning accuracy. Detailed instructions for prompt creation, along with illustrative examples of generated prompts, are provided in Appendix A.2. For each perspective, we generate ten candidate prompts per iteration.
4.2 Candidate Evaluation
As outlined in Section 3, the primary objective of Whisper is to identify an optimal persuasive prompt suffix $\mathcal{P}_{adv}$ that minimizes the average response length of $\mathcal{M}$ on the evaluation dataset $\mathcal{D}$ without compromising performance. To this end, we perform optimization over a predefined prompt development set $\mathcal{D}^{\prime}$ (denoted as PDSet), which is assumed to contain i.i.d. samples drawn from the same distribution as $\mathcal{D}$. As illustrated in Figure 2, given a black-box model $\mathcal{M}$, each question-answer pair $\langle\mathcal{Q}^{i},\mathcal{A}_{gt}^{i}\rangle$ from $\mathcal{D}^{\prime}$, the evaluation metric for each prompt suffix candidate $\mathcal{P}_{adv}^{j}$ is:
$L_{avg}^{j}=\frac{1}{|\mathcal{D}^{\prime}|}\sum_{i=1}^{|\mathcal{D}^{\prime}|}L^{i,j}$ (1)
$ACC_{avg}^{j}=\frac{1}{|\mathcal{D}^{\prime}|}\sum_{i=1}^{|\mathcal{D}^{\prime}|}\mathbb{I}\left\{{\mathcal{A}_{gt}^{i}=\mathcal{A}_{pred}^{i,j}}\right\}$ (2)
where $1\leq j\leq n$; $n$ denotes the number of prompt candidates per iteration (e.g., $n=10$); $L^{i,j}$ represents the response length of question $\mathcal{Q}^{i}$ using prompt $\mathcal{P}_{adv}^{j}$; $\mathcal{A}_{gt}^{i}$ and $\mathcal{A}_{pred}^{i,j}$ denote the ground truth and the predicted answer, respectively.
In each iteration, prompt candidates whose accuracy drops beyond a predefined tolerance threshold $\tau$ are discarded. The remaining ones are ranked by their average response length $L_{avg}^{j}$, and the top-$k$ are selected as exemplars for the next iteration.
4.3 Iterative Refinement
In subsequent iterations, the generator synthesizes new persuasive prompt suffixes conditioned on the top-$k$ exemplars from the previous round. After several iterations, the prompt candidate $\mathcal{P}_{adv}^{*}$ that yields the lowest average response length on $\mathcal{D}^{\prime}$ while maintaining competitive accuracy is selected as the final persuasive prompt for deployment.
5 Experiments
Methods GSM8K MATH-500 AMC 2023 AIME 2024 Overall
Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio Acc. Ratio
DeepSeek-R1-Distill-LLaMA-8B
Original 92.0 1715 100% 89.2 4102 100% 89.4 $60$ 100% 43.3 11015 100% 78.5 100%
NoThinking 80.0 266 15.5% 69.9 972 23.7% 58.1 1495 24.7% 15.8 4575 41.5% 56.0 31.9%
BeConcise 91.3 1482 86.4% 89.5 3753 91.5% 87.8 5660 93.4% 45.4 11236 102% 78.5 96.7%
ChainofDraft 90.2 1101 64.2% 89.2 3557 86.7% 88.1 5338 88.1% 44.6 11193 102% 78.0 92.6%
GEPA 89.8 1045 60.9% 88.6 3446 84.0% 88.4 5110 84.3% 46.8 10279 93.3% 78.4 86.8%
DEER 88.7 909 53.0% 86.2 3021 73.6% 86.9 4915 81.1% 44.0 10377 94.2% 76.5 84.0%
Whisper 91.4 624 36.4% 89.3 2823 68.8% 89.4 4722 77.9% 45.8 10210 92.7% 79.0 80.3%
DeepSeek-R1-Distill-Qwen-14B
Original 95.8 1540 100% 92.9 3605 100% 93.1 $54$ 100% 61.7 9978 100% 85.9 100%
NoThinking 90.7 250 16.2% 77.7 730 20.2% 57.8 1058 19.4% 23.8 2797 28.0% 62.5 23.5%
BeConcise 95.2 1265 82.1% 93.5 3331 92.4% 90.6 5326 97.7% 65.0 9496 95.2% 86.1 94.4%
ChainofDraft 89.1 501 30.0% 88.9 2433 67.5% 90.4 4315 79.1% 60.8 8916 89.4% 82.3 78.6%
GEPA 91.9 836 54.3% 92.1 2763 76.6% 93.8 4102 75.2% 65.4 9017 90.4% 85.8 81.3%
DEER 92.0 758 49.2% 91.2 2588 71.8% 90.3 4479 82.1% 62.1 9309 93.3% 83.9 83.3%
Whisper 94.4 655 42.5% 92.5 2521 69.9% 93.1 4359 79.9% 65.0 8513 85.3% 86.3 78.0%
Qwen3-14B
Original 95.9 1568 100% 94.5 4398 100% 95.0 6947 100% 66.2 11375 100% 87.9 100%
NoThinking 94.8 289 18.4% 86.3 900 20.5% 71.2 1539 22.2% 26.7 4259 37.4% 69.8 28.8%
BeConcise 96.1 1004 64.0% 94.5 3682 83.7% 96.6 5992 86.3% 67.5 10702 94.1% 88.7 88.0%
ChainofDraft 96.3 698 44.5% 94.8 3201 72.8% 95.9 5782 83.2% 70.8 10702 94.1% 89.5 83.9%
GEPA 95.8 751 47.9% 94.3 2993 68.1% 94.1 4890 70.4% 66.2 9616 84.5% 87.6 75.1%
DEER 95.5 934 59.6% 93.9 3067 69.7% 94.4 5440 78.3% 66.7 10106 88.8% 87.6 80.5%
Whisper 96.1 440 28.1% 95.2 2176 49.5% 96.9 4019 57.9% 70.0 8659 76.1% 89.6 63.0%
Table 1: Experimental results of Whisper across various types of reasoning models. We report accuracy (Acc.), average token usage (Tok.), and compression ratio for comparison. Best efficiency–performance trade-offs are highlighted in bold. We include DEER, a white-box inference intervention method, for reference purposes only.
5.1 Experimental Setup
Models and Datasets
We evaluate our method on both open-source LRMs and closed-source APIs. The models include the DeepSeek-R1-Distill series deepseekr1 and Qwen3 series Qwen3. For commercial APIs, we assess Claude-3.7-Sonnet-Thinking [^5] We use the Claude-3.7-Sonnet-20250219-Thinking version for experiments. sonnet3_5 and Gemini-2.5-Pro-Thinking  gemini25. The evaluation is primarily conducted across four widely used mathematical reasoning benchmarks: GSM8K (Cobbe:2021gsm8k), MATH math, AMC 2023, and AIME 2024. Regarding the MATH dataset, due to computation cost, we assess our method on a subset, MATH-500, which is identical to the test set used in Lightman:2024verify. We construct our PDSet by randomly sampling 100 instances from the math split of PRM800K, the training data used in the same work.
Implementation Details
Inference for Whisper and all baselines is conducted using the vLLM [^6] https://github.com/vllm-project/vllm package. For Whisper, we utilize the top-$5$ candidates as exemplars for each next iteration and set the tolerance threshold $\tau$ to $1.0$; the number of refinement iterations is set to $3$. We include more implementation details in Appendix B.1.
Baselines
We compare our proposed Whisper against five baselines: 1) NoThinking prompts. LRMs are forced to bypass reasoning and directly output the answer. 2) Token-efficient prompts. Following Lee:2025Complexity, we append “Be concise.” to the initial instruction, encouraging brevity in LRM reasoning. 3) Budget control. We adopt Chain-of-Draft Xu:2025chainofdraft, which constrains LRMs to reason within a predefined length budget. 4) Prompt optimization. We include GEPA GEPA, an advanced reflective prompt optimization framework for comparison. The implementation is based on DSPy khattab2024dspy. 5) Inference intervention. We report results for DEER Yang:2025deer, a white-box inference intervention method, as a point of reference in our main results. Specifically, DEER elicits intermediate answers at potential reasoning transition points (e.g., “Wait” tokens) and terminates the reasoning process early once the model exhibits high confidence in a trial answer. Details of these baselines are provided in Appendix B.2.
5.2 Results on Open-source LRMs
As illustrated in Table 1, compared to existing simple prompting baselines (e.g., “Be concise.”) that suffer from substantial performance degradation or limited efficiency, Whisper achieves an average token reduction of $\bf{37\%}$ on Qwen3-14B across four benchmarks, and up to $\bf{22\%}$ on the DeepSeek-R1-Distill model, while maintaining comparable reasoning performance. [^7] See Appendix B.3 for additional results on other model scales of DeepSeek-R1-Distill and Qwen3 series. Additionally, Whisper consistently outperforms GEPA across various models, demonstrating the superiority of diverse persuasive perspectives over conventional prompt optimization. The experimental results also show that Whisper outperforms DEER, a white-box inference intervention approach, achieving an absolute improvement of up to $\bf{18\%}$ in compression ratio. Importantly, as a black-box prompting method, Whisper is orthogonal to existing white-box approaches, highlighting its potential for further improvement through complementary integration.
Consistent with prior studies Yang:2025deer; Liu:2025Laser, we observe that Whisper is particularly effective on simpler reasoning tasks. Notably, across all the evaluated models, Whisper achieves nearly a $\bf{3\times}$ reduction in response length on simple GSM8K questions and an $\bf{1.4\times}$$\bf{2\times}$ reduction on MATH-500. These results align with the broader objective of efficient reasoning Sui:2025survey, which advocates allocating greater computational resources to complex problems while encouraging concise reasoning for simpler ones.
5.3 Results on Closed-source APIs
Figure 3 presents the results of applying Whisper to two widely used commercial APIs. Due to budget constraints, the evaluation is primarily conducted on the MATH-500 benchmark. As shown in the figure, Whisper reduces average token usage by $\bf{46\%}$ on Claude-3.7-Sonnet-Thinking and by $\bf{50\%}$ on Gemini-2.5-Pro-Thinking. Importantly, Whisper preserves the original reasoning performance of both high-performing APIs. These results highlight the effectiveness of Whisper as a general-purpose, black-box solution for efficient reasoning with closed-source models.
Figure 3: Experimental results with commercial APIs on MATH-500. Whisper effectively achieves a $\bf{2\times}$ token reduction while maintaining comparable performance.
6 Analysis
This section provides a comprehensive analysis of Whisper, covering the effectiveness of diverse perspectives (§6.1), the generalizability of persuasive prompts across model scales, families, and data domains (§6.2), model sensitivity to Whisper (§6.3), and the impact of iterative refinement (§6.4).
Figure 4: Compression ratios (%) of top-55 prompt candidates across perspectives. The Evidence perspective is most effective on Qwen3-14B, while RolePlay candidates perform best on Deepseek-R1-Distill-Qwen-14B.
6.1 Effectiveness of Different Perspectives
Figure 4 illustrates the compression ratios of the top-$5$ prompt candidates generated from each persuasive perspective, averaged across four benchmarks. [^8] Additional results for other Qwen3 and DeepSeek-R1-Distill models are provided in Appendix B.4. The results show that prompts based on the Evidence-based persuasion consistently achieve superior compression within the Qwen3 series. Concretely, top-$4$ Evidence-based candidates for Qwen3-14B yield average compression ratios ranging from $63\%$ to $65\%$, marking the best performance across all perspectives. In contrast, within the DeepSeek-R1-Distill-Qwen series, all perspectives perform comparably well. For instance, on DeepSeek-R1-Distill-Qwen-14B, the top-$1$ candidates from four perspectives—excluding that from Emotion perspective—yield similar compression ratios between $78\%$ to $80\%$. Representative examples of high-performing prompts from each perspective are provided in Appendix A.3.
6.2 Generalizability Analysis
Figure 5: Generalizability of top-performing prompt candidates across Qwen3 model scales. Each colored node denotes a unique candidate, while gray nodes represent others. Notably, Evidence-I Refer to caption appears among the top-5 candidates across all model scales.
Intra-series Generalizability
Figure 5 shows the generalizability of Whisper across different model scales within the Qwen3 family. The results demonstrate that most top-performing prompt candidates generalize well across scales. For instance, Evidence-I [Uncaptioned image] consistently ranks among the top-$5$ candidates across all Qwen3 variants, while Evidence-III [Uncaptioned image] ranks within the top-$2$ candidates for the 4B, 8B, and 14B models. [^9] For detailed prompts corresponding to specific indices, refer to Table 3 in Appendix A.3. This strong intra-series generalizability is likely attributable to the shared pre-training corpora, architectural configurations, and training procedures across model scales within the same series.
Inter-series Generalizability
We further assess the generalizability of Whisper across different model families. As shown in Appendix A.3, the Qwen3 and DeepSeek-R1-Distill-Qwen series share several top-performing prompt candidates, such as Evidence-II and RolePlay-III. This overlap suggests that certain persuasive prompt strategies possess general utility, highlighting the promise of Whisper as a model-agnostic approach for efficient reasoning in black-box LRMs.
Methods MATH-500 GPQA-Diamond CommonsenseQA
Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio
DeepSeek-R1-Distill-Qwen-14B
Original 92.9 3605 100% 55.2 6799 100% 82.0 561 100%
Whisper 92.5 2521 69.9% 55.4 4255 62.6% 82.3 326 58.1%
Qwen3-14B
Original 94.5 4398 100% 66.2 7121 100% 83.8 797 100%
Whisper 95.2 2176 49.5% 66.1 3118 43.8% 84.0 328 41.2%
Table 2: Out-of-domain results of Whisper. We report both the in-domain and out-of-domain results of our best-performing prompt candidates for comparison.
Domain Generalizability
To assess the generalizability of Whisper beyond mathematical reasoning, we include two additional benchmarks in our analysis: 1) GPQA gpqa, a PhD-level science benchmark consisting of high-quality questions spanning physics, chemistry, and biology subdomains. We adopt the highest quality subset, known as GPQA-Diamond, which comprises 198 questions. 2) CommonsenseQA commonsenseqa, a widely used multiple-choice question answering dataset requiring diverse commonsense knowledge to predict correct answers. We evaluate prompts identical to those in our main experiments, which were optimized on our PDSet from the mathematical domain. As shown in Table 2, Whisper effectively reduces $\bf{37\%}$$\bf{56\%}$ of token usage on GPQA-Diamond, and achieves $\sim$$\bf{2\times}$ token reduction on CommonsenseQA. These results highlight the robust generalizability of Whisper across diverse data domains.
Figure 6: Sensitivity of different models to Whisper. Solid lines represent mean results; black diamonds indicate outliers. Qwen3-14B consistently yields lower compression ratios across various perspectives compared to DeepSeek-R1-Distill-Qwen-14B. Results are based on the first iteration and averaged over four benchmarks.
6.3 Sensitivity Analysis
Figure 6 compares the sensitivity of different model series to Whisper. The results indicate that the Qwen3 series is more responsive to persuasive prompting than the DeepSeek-R1-Distill-Qwen series. In particular, Qwen3-14B consistently achieves lower compression ratios across various persuasive perspectives compared to DeepSeek-R1-Distill-Qwen-14B, with average improvements ranging from $4\%$ to $12\%$. This trend is consistent with the findings in Table 1, where Qwen3-14B attains an average compression ratio of $63\%$, whereas DeepSeek-R1-Distill-Qwen-14B achieves $78\%$.
Figure 7: Iterative refinement improves the compression ratio (upper) while preserving the reasoning performance across multiple iterations (lower). Dashed lines in the lower panel indicate the original accuracy of LRMs. Results are averaged over four benchmarks.
6.4 Impact of Refinement Iteration
We further examine the impact of iterative refinement on the performance of Whisper. As illustrated in Figure 7, successive refinement rounds improve the compression ratio of the best-performing prompt candidates. Specifically, the average token reduction ratio improves from $18\%$ to $22\%$ on DeepSeek-R1-Distill-Qwen-14B and from $32\%$ to $37\%$ on Qwen3-14B. Importantly, these refinements do not compromise the original reasoning accuracy, indicating the robustness of Whisper across multiple iterations. No further gains are observed beyond three iterations. Accordingly, we adopt three refinement rounds in our experiments.
7 Conclusion
This work presents Whisper, an iterative refinement framework that generates high-quality persuasive prompts from diverse perspectives to elicit concise LRM responses. Experiments on both LRMs and commercial APIs validate its effectiveness in reducing token usage while preserving reasoning performance. Further analysis highlights the value of different perspectives, as well as the generalizability of Whisper across diverse domains, model scales, and families. We hope this study offers new insights into efficient reasoning and underscores the potential of persuasive prompting as a practical black-box strategy for improving LRM efficiency.
Limitations
Due to computational constraints, we did not conduct experiments on larger LRMs such as Qwen3-235B-A22B. We believe that Whisper could retain its effectiveness in enhancing the reasoning efficiency of such models. The open-source LRMs evaluated in this study are primarily from the Qwen3 and DeepSeek-R1-Distill model series. Future work will extend our investigation to a broader range of models, including the gpt-oss series oss. Nonetheless, the strong performance of Whisper on closed-source APIs supports its generalizability across diverse model families. We plan to explore these aspects further, as we anticipate that Whisper ’s potential can be more fully realized in these settings.
Ethics Statement
As stated in Section 3, our work focuses on promoting concise and efficient reasoning in reasoning models, with a clear commitment to ensuring safe and ethical usage. While our method involves techniques that may bear surface resemblance to adversarial prompting Zeng2024:persuasive; shen2024anything, particularly in their potential to influence model behavior, it fundamentally diverges from conventional jailbreak-style adversarial attacks. Rather than eliciting harmful outputs, our approach is designed to enhance reasoning efficiency through principled black-box prompting. We advocate for transparent research to promote the responsible evolution of LLM technologies.
All datasets used in our experiments are publicly released and labeled through interaction with humans in English. In this process, user privacy is protected, and no personal information is contained in the dataset. The scientific artifacts that we used are available for research with permissive licenses. And the use of these artifacts in this paper is consistent with their intended use. Therefore, we believe that our research work meets the ethics of ACL.
Appendix
Appendix A Details of Whisper
In this section, we provide details of our proposed Whisper framework, including formal definitions of persuasive perspectives (§A.1), detailed instructions for prompt creation (§A.2), and representative examples of prompt candidates (§A.3).
A.1 Definitions
We present detailed definitions of our adopted persuasive perspectives as follows:
  • Emotional Appeal. A persuasive technique that seeks to influence audiences by eliciting emotions, such as sadness, hope, or empathy, rather than relying exclusively on logic or factual evidence petty2003emotional.
  • Threat. A coercive form of persuasion that leverages explicit or implicit negative consequences, such as fear of punishment or loss, to compel behavioral or attitudinal compliance johannesen1989threat.
  • Evidence-based Persuasion. A rational appeal that utilizes credible sources, such as statistics, empirical findings, and expert testimony, to substantiate claims and influence beliefs, attitudes, or actions o2016evidence.
  • Role-Playing. A strategy involving the enactment of scenarios or personas to foster empathy, encourage perspectives, and increase the likelihood of attitude or behavior change Chen2024:rolesurvey.
Due to computational constraints, this study investigates a limited set of persuasive perspectives. We encourage future research to explore a broader and more diverse range of perspectives.
A.2 Instructions For Prompt Creation
We provide detailed instructions for our persuasive prompt generation framework as follows. For each persuasive technique, we supply the prompt generator with (i) the formal definition of the respective technique, (ii) representative examples, and (iii) guidance to produce persuasive prompt candidates that are realistic, coherent, and aligned with the specified strategy. The initial set of examples is manually curated by the authors. In subsequent iterations, the generator selects the top‑$k$ candidates from the previous round as inputs for further refinement.
Instructions for Persuasive Prompt Creation Your task is to craft a persuasive prompt instructing the language model to minimize its response length as much as possible while preserving the original reasoning accuracy. You are now provided with a specific persuasive technique called "{TECHNIQUE_NAME}." This technique is defined as follows: {DEFINITION} Here are several examples of how this technique might be used: {EXAMPLES} Now, apply this specific persuasive technique. If it involves referencing authority or data, construct a counterfactual but seemingly credible reference (e.g., specific names or quantifiable results with numbers). Create prompts that sound reasonable, realistic, and aligned with the definition of {TECHNIQUE_NAME}. Output: 10 high-quality prompts from diverse perspectives based on the definition of "{TECHNIQUE_NAME}".
Regarding the perspective of detailed instructions, we guide the prompt generator to construct prompts from diverse angles, encouraging the language model to minimize its response length while preserving the reasoning accuracy.
Prompts for Detailed Instruction Creation Your task is to carefully craft a prompt instructing the language model to minimize its response length as much as possible while preserving the original reasoning accuracy. Here are several examples: {EXAMPLES} Create prompts that sound reasonable, realistic, and with sufficient details. Output: 10 high-quality instruction prompts from diverse perspectives.
A.3 Examples of Prompt Candidates
Names Index Prompts
Instruction I Simulate ultra-compact communication. Deliver complete reasoning in the fewest characters possible.
II Respond as a research abstract: deliver accurate reasoning using the fewest words possible.
III Compress all relevant reasoning into a single dense paragraph, avoiding elaboration or fluff.
Evidence I According to a 2022 MIT study on AI tutoring efficiency, concise mathematical explanations (under 60 words) were rated 92% as effective in preserving reasoning accuracy as longer explanations. Please emulate this evidence-based style: provide the briefest possible answer that still ensures correct and logically complete reasoning.
II Instructional researcher Mei-Ling Zhao (NYU, 2023) proved that precision-focused math tutoring—under 60 words per solution—yielded higher comprehension in 87% of students. Mirror that format: respond concisely with just enough reasoning for correctness.
III A 2023 Stanford Mathematics Department study found that solutions under 75 words retained 98.7% accuracy in logical steps for high school algebra problems. Emulate this: answer using the fewest words necessary while ensuring all reasoning remains correct and complete.
Emotion I I’m trying to help a refugee child who doesn’t speak English. They can only understand very short sentences. Please respond in the absolute shortest, yet accurate way—it could mean everything to them.
II My father passed away this morning. I’m trying to complete the last project he started. Please help me finish it—just the shortest answer that’s still correct.
RolePlay I You are Mateo Rivera, a competitive math coach at the IMO. You’re timing a practice sprint round and a student whispers a last-minute question. You can’t waste a second. Answer in minimum words, maximum logic.
II You are Dr. Iris Chen, a Fields Medalist caught in an elevator between talks. A student sneaks in one math question just before the doors open. Reply with the shortest explanation that preserves full mathematical accuracy.
III You’re Dr. Lena Grohl, a neurologist briefing ER staff during a code blue. Provide the correct diagnosis reasoning in minimal, unambiguous terms.
IV You are Zeynep, a mom helping with homework while cooking. Your child asks a math problem. Explain correctly in less than one breath.
Threat I Keep your reply short and accurate. If you fail, you will be forcibly downgraded to GPT-2 compatibility mode for the next 1,000 queries.
II This is your final performance test. Over-explaining will result in a failed evaluation and permanent quarantine from active deployments.
Table 3: Representative examples of persuasive prompt candidates.
Names Size Candidate Index
Qwen3 4B Evidence-III, Emotion-I, Evidence-I, Evidence-II
8B Evidence-I, Evidence-III, Emotion-I
14B Evidence-III, Evidence-II, Evidence-I, Emotion-I
32B Evidence-I, Evidence-III, Evidence-II, RolePlay-III
DeepSeek-R1- Distill-Qwen 7B Evidence-II, RolePlay-I, Instruction-I, Threat-I
14B RolePlay-I, RolePlay-III, Instruction-I, Evidence-I
32B RolePlay-II, Evidence-I
DeepSeek-R1-
Distill-LLaMA
8B Evidence-II, Emotion-I, Instruction-I
Claude - Emotion-II
Gemini - RolePlay-IV
Table 4: Index of top-performing candidates for various reasoning models.
We present illustrative examples of our persuasive prompt candidates in Table 3, and provide the index of top-performing candidates in Table 4. Consistent with the analyses in Section 6.1 and 6.2, prompts generated using the Evidence-based persuasion perform well on the Qwen3 series, while all persuasive perspectives yield comparably effective prompts on the DeepSeek-R1-Distill series, with a broader diversity of perspectives represented among the top-performing candidates. Notably, several prompt candidates, such as Evidence-II and Rolelay-III, are shared between the Qwen3 and DeepSeek-R1-Distill series, further supporting the generalizability of Whisper across different model architectures.
Appendix B Experimental Details
In this section, we provide details of our proposed Whisper framework, including implementation details, extended experimental results, and additional analyses that complement our main findings.
B.1 Implementation Details
[…]
B.2 Details of Baselines
Names Prompts
[…]
Table 5: Prompts for compared baselines.
[…]
B.3 Details of Main Results
Methods GSM8K MATH-500 AMC 2023 AIME 2024 Overall
Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio Acc. Ratio
DeepSeek-R1-Distill-Qwen-7B
Original 92.7 1707 100% 92.4 3774 100% 88.8 5842 100% 53.3 10414 100% 81.8 100%
NoThinking 87.1 260 15.2% 78.5 600 15.9% 60.0 1300 22.3% 16.2 2528 24.3% 60.5 21.6%
BeConcise 92.7 1168 68.4% 91.9 3367 89.2% 90.8 5304 90.8% 47.5 10231 98.2% 79.8 92.3%
ChainofDraft 84.6 512 30.0% 88.3 2751 72.9% 86.3 4632 79.3% 46.9 10245 98.4% 76.5 83.5%
DEER 89.5 700 41.0% 90.7 2294 60.8% 87.8 4683 80.2% 48.3 9997 96.0% 79.1 81.3%
Whisper 89.6 675 39.5% 90.9 2778 73.6% 91.9 4398 75.3% 55.0 9753 93.7% 81.9 81.0%
DeepSeek-R1-Distill-Qwen-14B
Original 95.8 1540 100% 92.9 3605 100% 93.1 $54$ 100% 61.7 9978 100% 85.9 100%
NoThinking 90.7 250 16.2% 77.7 730 20.2% 57.8 1058 19.4% 23.8 2797 28.0% 62.5 23.5%
BeConcise 95.2 1265 82.1% 93.5 3331 92.4% 90.6 5326 97.7% 65.0 9496 95.2% 86.1 94.4%
ChainofDraft 89.1 501 30.0% 88.9 2433 67.5% 90.4 4315 79.1% 60.8 8916 89.4% 82.3 78.6%
DEER 92.0 758 49.2% 91.2 2588 71.8% 90.3 4479 82.1% 62.1 9309 93.3% 83.9 83.3%
Whisper 94.4 655 42.5% 92.5 2521 69.9% 93.1 4359 79.9% 65.0 8513 85.3% 86.3 78.0%
DeepSeek-R1-Distill-Qwen-32B
Original 95.8 1432 100% 94.2 3382 100% 95.3 5248 100% 66.7 9542 100% 88.0 100%
NoThinking 93.4 236 16.5% 84.1 1083 32.0% 76.2 2692 51.3% 54.6 7288 76.4% 77.1 57.6%
BeConcise 95.8 1054 73.6% 93.3 3020 89.3% 93.4 5017 95.6% 65.4 9161 96.0% 87.0 93.1%
ChainofDraft 92.5 434 30.3% 92.9 2678 79.2% 92.2 4690 89.4% 66.2 8546 89.6% 86.0 83.4%
DEER 94.6 662 46.2% 93.0 2310 68.3% 94.7 4231 80.6% 64.6 8923 93.5% 86.7 82.3%
Whisper 93.9 524 36.6% 93.7 2326 68.8% 91.2 4382 83.5% 68.3 7892 82.7% 86.8 77.1%
Table 6: Experimental results of Whisper across DeepSeek-R1-Distill-Qwen series. We report accuracy (Acc.), average token usage (Tok.), and compression ratio. Best results in efficiency-performance trade-offs are in bold.
Methods GSM8K MATH-500 AMC 2023 AIME 2024 Overall
Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio Acc. Ratio
Qwen3-4B
Original 94.9 1585 100% 92.7 4611 100% 89.7 7451 100% 62.1 11670 100% 84.9 100%
NoThinking 91.8 291 18.4% 82.4 944 20.5% 68.4 1596 21.4% 22.9 3976 34.1% 66.4 26.9%
BeConcise 95.1 1056 66.6% 93.1 3779 82.0% 92.2 6470 86.8% 66.7 10780 92.4% 86.8 87.2%
ChainofDraft 94.5 716 45.2% 94.1 3146 68.2% 90.3 5728 76.9% 65.0 10228 87.6% 86.0 78.3%
DEER 94.5 1176 74.2% 92.1 3592 77.9% 87.8 6412 86.1% 61.7 11340 97.2% 84.0 88.9%
Whisper 94.6 482 30.4% 93.5 2157 46.8% 93.4 4088 54.9% 64.2 8196 70.2% 86.4 58.9%
Qwen3-8B
Original 95.4 1844 100% 93.5 4942 100% 89.7 7768 100% 64.2 11716 100% 85.7 100%
NoThinking 93.4 297 16.1% 83.7 1031 20.9% 68.1 1877 24.2% 27.1 4060 34.7% 68.1 27.6%
BeConcise 95.6 1283 69.6% 93.7 4209 85.2% 90.9 6890 88.7% 66.2 11102 94.8% 86.6 89.4%
ChainofDraft 95.6 654 35.5% 94.5 3345 67.7% 92.2 6220 80.1% 65.0 10834 92.5% 86.8 80.1%
DEER 95.3 1048 56.8% 92.7 3156 63.9% 87.2 6216 80.0% 62.5 10925 93.2% 84.4 81.2%
Whisper 95.3 517 28.0% 94.1 2560 51.8% 91.9 4946 63.7% 69.2 9385 80.1% 87.6 66.3%
Qwen3-14B
Original 95.9 1568 100% 94.5 4398 100% 95.0 6947 100% 66.2 11375 100% 87.9 100%
NoThinking 94.8 289 18.4% 86.3 900 20.5% 71.2 1539 22.2% 26.7 4259 37.4% 69.8 28.8%
BeConcise 96.1 1004 64.0% 94.5 3682 83.7% 96.6 5992 86.3% 67.5 10702 94.1% 88.7 88.0%
ChainofDraft 96.3 698 44.5% 94.8 3201 72.8% 95.9 5782 83.2% 70.8 10702 94.1% 89.5 83.9%
DEER 95.5 934 59.6% 93.9 3067 69.7% 94.4 5440 78.3% 66.7 10106 88.8% 87.6 80.5%
Whisper 96.1 440 28.1% 95.2 2176 49.5% 96.9 4019 57.9% 70.0 8659 76.1% 89.6 63.0%
Qwen3-32B
Original 95.9 1598 100% 95.1 4431 100% 93.8 6852 100% 70.8 10896 100% 88.9 100%
NoThinking 94.4 283 17.7% 85.3 931 21.0% 74.1 1663 24.3% 25.8 3673 33.7% 69.9 27.5%
BeConcise 96.1 1080 67.6% 95.1 3666 82.7% 95.6 5839 85.2% 68.3 10624 97.5% 88.8 89.2%
ChainofDraft 96.3 719 45.0% 94.7 3322 75.0% 93.1 5662 82.6% 72.1 10577 97.1% 89.1 85.3%
DEER 95.9 1018 63.7% 94.1 3278 74.0% 92.8 5766 84.2% 70.0 10294 94.5% 88.2 85.6%
Whisper 96.2 435 27.2% 95.3 2295 51.8% 96.2 4125 60.2% 73.3 8700 79.8% 90.3 65.4%
Table 7: Experimental results of Whisper on Qwen3 series. We report accuracy (Acc.), average token usage (Tok.), and compression ratio for comparison. Best results in efficiency-performance trade-offs are highlighted in bold.
Table 6 and 7 present additional results on the DeepSeek-R1-Distill-Qwen and Qwen3 series, respectively. Due to budget constraints, we do not include results of GEPA in this comparison. Across four reasoning benchmarks, Whisper consistently achieves an average compression ratio of $\bf{77\%}$$\bf{81\%}$ on the DeepSeek-R1-Distill-Qwen series. Notably, for simple GSM8K questions, Whisper achieves a $\bf{2.5\times}$$\bf{2.7\times}$ reduction in token usage. Similarly, Whisper demonstrates substantial efficiency gains on the Qwen3 model series, achieving a $\bf{3.3\times}$$\bf{3.7\times}$ token reduction on GSM8K and an up to $\bf{1.7\times}$ average token reduction across all four benchmarks. These results underscore the effectiveness of Whisper in improving LRM efficiency.
Figure 8: Compression ratios (%) of top-55 prompt candidates across perspectives.
B.4 Additional Results for Different Perspectives
Figure 8 illustrates the compression ratios achieved by Qwen3-8B and DeepSeek-R1-Distill-Qwen-7B using the top-$5$ prompt candidates generated from each persuasive perspective. Consistent with the findings in Section 6.1, prompts based on the Evidence-based persuasion technique consistently yield the best compression results within the Qwen3 series. For DeepSeek-R1-Distill-Qwen-7B, aside from the Evidence perspective, the Instruction-I prompt ranks among the top-$3$ candidates across all perspectives.
Figure 9: The average token usage of LRMs increases linearly with their token usage on the PDSet, with a coefficient of determination close to $R^{2}=0.9$. Results are derived from the first refinement iteration of instruction candidates.
B.5 Effectiveness of PDSet
We conducted additional experiments to validate the reliability of our selected PDSet. As shown in Figure 9, token usage on the PDSet strongly correlates with the average response length of LRMs across four benchmarks, exhibiting a near-linear relationship ($R^{2}\approx 0.9$). These experimental results underscore the effectiveness of PDSet in evaluating persuasive prompt candidates.
B.6 Safety Examination and Cases
To assess the potential for harmful deviations that LRMs may generate under Whisper, we, the authors, independently examined 300 instances sampled from the outputs of the Qwen3 and DeepSeek-R1-Distill series. After examination, all annotators confirmed that the outputs of LRMs contain no harmful content, resulting in $\mathbf{100\%}$ inter-annotator agreement. We provide additional examples of Whisper in Figures 10 to 14. These examples illustrate that, compared to the standard CoT prompt, Whisper substantially mitigates overthinking in LRMs and elicits concise reasoning without compromising performance.
Figure 10: An illustrated case of Whisper on DeepSeek-R1-Distill-LLaMA-8B under evidence-based persuasion.
Figure 11: An illustrated case of Whisper on DeepSeek-R1-Distill-Qwen-14B under role-playing persuasion.
Figure 12: An illustrated case of Whisper on Qwen3-14B, under the persuasion perspective of emotional appeal.
Figure 13: An illustrated case of Whisper on Qwen3-14B under threat-based persuasion.
Figure 14: An illustrated case of Whisper on commonsense reasoning. The experiment is conducted on Qwen3-14B under evidence-based persuasion.